what is the molecular formula of the structure?
how many sigma bonds are present in the steucture as drawn?
how many pi bonds are present in the structure as draw?
what is the hybridization of every carbon in the left most ring

IN
Answer #1

O I Il H 2 X 307 V T + н на mol for = Cq HQ ONSB, Sigma bonds = 21 (as drawn => 13) pi bonds = 3 (as drawn 3) C = 56² C = sp²